OpenStudy (anonymous):

What is the fourth term in this multiplication pattern? multiplication pattern using the formula 2 • 6n – 1 A. 12 B. 72 C. 432 D. 2592

OpenStudy (anonymous):

@crazysingh

OpenStudy (campbell_st):

well substiute n = 1, then n= 2 then n = 3 etc and evalaute

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

im in 6th grade

OpenStudy (campbell_st):

and is the question actually \[2 \times 6^{n - 1}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@crazysingh

OpenStudy (campbell_st):

on so here is n = 1 \[2 \times 6^{ 1 - 1} = 2 \times 6^{0}\] just find the value of the equation above then multiply it by 6 to get a new answer repeat the process twice more for your answer...

OpenStudy (anonymous):

72

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

216

OpenStudy (campbell_st):

so you need 2 x 6 x 6 x 6 =

OpenStudy (anonymous):

432

OpenStudy (anonymous):

so its C. thx so much!!!

OpenStudy (anonymous):

Glenn invested $18,000 at 5% interest compounded annually. How much interest will Glenn earn in 3 years? A. $20,837.25 B. $19,826.24 C. $2837.25 D. $2340.23

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

wow... I didn't think compound interest was a 6th grade problem...

OpenStudy (anonymous):

it is

OpenStudy (campbell_st):

lol...ok... then do you know the formula...

OpenStudy (anonymous):

no

OpenStudy (campbell_st):

ok... well here is a link https://qrc.depaul.edu/StudyGuide2009/Notes/Savings%20Accounts/Compound%20Interest.htm I think it explains it well... good luck

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ r=rate\qquad t=years\qquad n=periods\qquad P=principal\)

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

this is hard!

OpenStudy (anonymous):

Glenn invested $18,000 at 5% interest compounded annually. How much interest will Glenn earn in 3 years? A. $20,837.25 B. $19,826.24 C. $2837.25 D. $2340.23

OpenStudy (anonymous):

is the question

OpenStudy (anonymous):

*sigh*

OpenStudy (anonymous):

oh yes i knew the formula i just didnt get what u wre saying

OpenStudy (campbell_st):

the only different your n value = 1

OpenStudy (jdoe0001):

P= principal or original amount = 18,000 r = rate in decimal format 3% = 0.03 t = years n = periods per year, well is annually, thus once a year, thus period = 1 thus \(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ r=0.03\qquad t=3\qquad n=1\qquad P=18,000 \\ \quad \\ A=18,000\left(1+\frac{0.03}{1}\right)^{1\cdot 3}\)

OpenStudy (anonymous):

umm hard so hard!!!

OpenStudy (jdoe0001):

thus campbell_st 's comment -> wow... I didn't think compound interest was a 6th grade problem...

OpenStudy (campbell_st):

here is a link to an online calculator that will help you http://web2.0calc.com/

OpenStudy (anonymous):

huh???!!!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do i do though on the calculator??!!

OpenStudy (anonymous):

b

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

plug in your values I gather

OpenStudy (anonymous):

whatever im leaving cya ugh im gonna get a F!

OpenStudy (anonymous):

im dum!

OpenStudy (anonymous):

stupid

OpenStudy (anonymous):

i hate myself!!!

OpenStudy (anonymous):

BYE!

OpenStudy (jdoe0001):

keep in mind that -> \(\bf A=18,000\left(1+\frac{0.03}{1}\right)^{1\cdot 3}\implies A=18,000\left(1+0.03\right)^3 \\ \quad \\ \implies A=18,000\left(1.03\right)^3\) you know maybe as campbell_st said, it may just be too premature of a chapter, which would not be your fault really assuming the section on compound interests hasn't been covered yet

OpenStudy (anonymous):

give me the answer plz!!!

OpenStudy (anonymous):

b im just putting b

rvc (rvc):

angel u r in which degree or say class

OpenStudy (anonymous):

if i get under an 80% its fault

OpenStudy (anonymous):

im in 6th grade!

rvc (rvc):

usa education

OpenStudy (anonymous):

i got a 40% WHAT THE HECK THAT feather HE GOT ME WRONG!

OpenStudy (anonymous):

i said that b it ch not feather!

OpenStudy (anonymous):

cya!

rvc (rvc):

woo cool! good keep going